# Language Models as Causal Effect Generators

### v1

@@ Abstract
We present a framework for large language model (LLM) based data generation with controllable causal structure. In particular, we define a procedure for turning any language model and any directed acyclic graph (DAG) into a sequence-driven structural causal model (SD-SCM). Broadly speaking, an SD-SCM is a causal model with user-defined structure and LLM-defined structural equations. We characterize how an SD-SCM allows sampling from observational, interventional, and counterfactual distributions according to the desired causal structure. We then leverage this procedure to propose a new type of benchmark for causal inference methods, generating individual-level counterfactual data without needing to manually specify functional relationships between variables. We create an example benchmark consisting of thousands of datasets, and test a suite of popular estimation methods on these datasets for average, conditional average, and individual treatment effect estimation, both with and without hidden confounding. Apart from generating data, the same procedure also allows us to test for the presence of a causal effect that might be encoded in an LLM. This procedure can underpin auditing LLMs for misinformation, discrimination, or otherwise undesirable behavior. We believe SD-SCMs can serve as a useful tool in any application that would benefit from sequential data with controllable causal structure.
@@ 1 Introduction
All causal inference methods have to deal with the same fundamental underlying problem: we can only observe one state of the world at a given moment in time. If we wish to reason about what caused what, we — in some way or other — will have to reason about a hypothetical world (or worlds) that we didn’t observe. This issue is often referred to more formally as The Fundamental Problem of Causal Inference [28]: a given individual can only receive one treatment at a time, so we can only ever observe one of their potential outcomes (or counterfactuals). Overcoming this fundamental problem not only makes causal inference a challenge to perform, but also a challenge to evaluate with realistic data. Our main focus — data generation with known causal structure — is thus an important tool both for the development and the evaluation of methods that deal with causality.
A well-established strategy for causal effect estimation uses randomization of a treatment to estimate average treatment effects (ATEs) [50], capturing the effect of a treatment on average across many individuals. Because ATEs are feasible to isolate with proper study design, many realistic datasets exist with ‘true’ or at least widely-accepted ATEs [38, 15, 30]. If, instead, inferences about individual treatment effects (ITEs) are desired, making additional modeling assumptions can allow us to reason about counterfactuals at an individual level; unfortunately, so long as counterfactuals remain unobservable, these assumptions remain untestable. This means that there is no easy way to collect realistic data for ITEs. [^1] It is common in literature on individual-level causal inference to draw no distinction between ITEs and conditional average treatment effects (CATEs) [55], but we emphasize these two quantities as distinct: the CATE is the conditional expectation of the ITE, which typically will not explain all ITE variation [39]. In particular, the counterfactual outcome $\tilde{y}$ and/or treatment assignment $\tilde{t}$ must always be manually generated, even if covariates $\tilde{x}$ are based on real data (see, e.g., [42]). Our method is aimed at using LLMs to provide this type of data in a flexible manner.
@@
Figure 1: (a) A sequence-driven structural causal model (SD-SCM) uses a language model to sample data according to any desired causal relationships between variables (blue solid arrows represent an example set of possible DAGs). Any variables whose values are sampled from the language model will potentially share the language model as a common cause (red dashed arrows), unless sampled manually, e.g., uniformly. (b) However, the effects of red dashed arrows are not necessarily significant. For example, the variable weather is a strong confounder (top) or strong collider (bottom), depending on which DAG is chosen. All structural equations are determined implicitly by whatever has already been encoded in the language model (see Section 3.3 for the full example).
Contribution.
We define a procedure for turning any language model and any DAG into a sequence-driven structural causal model (SD-SCM). We characterize in Section 3 how an SD-SCM provides access to observational, interventional, and counterfactual distributions over sequential data according to the desired causal structure, where structural equations are defined implicitly via language model (see Figure 1). In Section 4, we use SD-SCMs to create an example benchmark for causal effect estimation and test a suite of popular estimation methods across ATE, CATE, and ITE estimation. Section 5 demonstrates how this same technique can underpin the auditing of language models for (un)desirable causal effects. All benchmark datasets as well as code for SD-SCM-based data generation is available on GitHub. [^2] https://github.com/lbynum/sequence-driven-scms We believe SD-SCMs can be used across many applications as a general tool for imposing causal structure on sequentially generated data.
@@ 2 Related work
[12] lay out four categories of commonly-used methods for semi-synthetic data generation with known causal effects: (1) simulating treatment effects using real baseline outcomes [e.g., 33]; (2) using real covariates but simulating response surfaces [e.g., 58, 20, 27]; (3) performing biased sampling of randomized data [e.g., 24, 15]; and (4) constructing (proxies of) counterfactuals and interventions from real or empirical data [e.g., 42, 23]. The paradigm of fitting models to real data and then sampling synthetic data from the fit models is common in many works [52, 45]. In this area, the most closely related works to ours in spirit are those that fit generative models to real datasets such that treatments, outcomes, and covariates — in effect, entirely new datasets — can be sampled, such as [2] and [45]. While such methods are similar in that they rely on generative models, they are fundamentally different from ours, as they are based on individual datasets that already exist (and already have a fixed causal structure), rather than allowing for arbitrary causal structures to be imagined by a user and then parameterized by a generative model. Our setup is akin to a high-fidelity simulation environment [e.g., 44] that provides empirical counterfactual data [23], but without needing to manually design all aspects of the simulation, and in a manner that is instead based on natural language. This work is also loosely related to methods that parameterize structural causal models (SCMs) with generative models or other deep learning components, such as [47, 51], but such methods are geared towards counterfactual inference and learning causal relationships from existing data, rather than flexible data generation.
@@
Our work is not the first to suggest that language models can generate outputs that have casual structure. Many works aim to augment language models with the ability to generate counterfactual data [7, 40, 6, 25, 22]. Counterfactuals and causal reasoning are useful across various natural language processing (NLP) tasks, making this capability of particular interest for ongoing LLM research [57], and language models with causal reasoning capabilities have a wide variety of applications both within and beyond NLP [54, 31, 59, 41, 18, 37, 32, 22]. We are also not the first to point out that counterfactual data generation with language models is useful for understanding the internal ‘world model’ constructed by an LLM and auditing for bias [7, 21]. However, our focus in this work is fundamentally different: rather than explore what language models can do with causal reasoning and causal structure across different types of tasks, we instead focus on how to generate data given a desired causal structure. This capability has important use-cases for downstream tasks like the one we demonstrate here — generating treatment effects to benchmark effect estimation methods. But more broadly, we provide a generalization of how sequence data and structural causal models can be combined in order to flexibly generate observational, interventional, and counterfactual data for whatever purpose it might be useful.

### v2

@@ Abstract
In this work, we present sequence-driven structural causal models (SD-SCMs), a framework for specifying causal models with user-defined structure and language-model-defined mechanisms. We characterize how an SD-SCM enables sampling from observational, interventional, and counterfactual distributions according to the desired causal structure. We then leverage this procedure to propose a new type of benchmark for causal inference methods, generating individual-level counterfactual data to test treatment effect estimation. We create an example benchmark consisting of thousands of datasets, and test a suite of popular estimation methods for average, conditional average, and individual treatment effect estimation. We find under this benchmark that (1) causal methods outperform non-causal methods and that (2) even state-of-the-art methods struggle with individualized effect estimation, suggesting this benchmark captures some inherent difficulties in causal estimation. Apart from generating data, this same technique can underpin the auditing of language models for (un)desirable causal effects, such as misinformation or discrimination. We believe SD-SCMs can serve as a useful tool in any application that would benefit from sequential data with controllable causal structure.
Language Models as Causal Effect Generators
Lucius E.J. Bynum New York University lucius@nyu.edu          Kyunghyun Cho New York University kyunghyun.cho@nyu.edu
@@ 1 Introduction
Figure 1: Illustration of a sequence-driven structural causal model (SD-SCM), which uses a language model to sample data according to a user-specified DAG. Any variables whose values are sampled from the language model will potentially share the language model as a common cause (dashed arrows), unless sampled manually, e.g., uniformly.
Reasoning about counterfactuals plays a fundamental role in understanding cause and effect, both in theory and in practice. Unfortunately, counterfactuals are also fundamentally unobservable (Holland, 1985) and must always be simulated. In this work, we leverage language models (LMs) to help simulate counterfactual data in a user-controlled manner. To achieve this, we borrow the conditional distributions of a pre-trained LM in order to parameterize a structural causal model, based on an input directed acyclic graph (DAG) over variables expressed in natural language. This procedure allows us to simulate true counterfactual data — to observe both potential outcomes — but, crucially, without manually specifying functional relationships between variables. Instead, the specification of structural equations becomes data-driven. We explore how this data-driven approach can enable the specification of causal models for complex settings with less reliance on human expertise or creativity to manually specify relationships between variables.
Figure 2: Toy example showing SD-SCMs that use GPT-2 (Radford et al., 2019) to generate observational and counterfactual data corresponding to user-specified DAGs. In one case, the red node (weather, $\tilde{w}$) is a confounder. In the other case, $\tilde{w}$ is a collider. Plots on the right show that despite possible effects from dashed arrows, and DAGs that may contradict what we expect to happen in the real world, the generated data are indeed consistent with $\tilde{w}$ as a confounder or a collider.
Many use-cases are possible for sequential data (like text) with controllable causal structure. The main use-case we explore in this work is the development of a new type of benchmark for causal inference — a benchmark for conditional average and individual treatment effect estimation, where neither the counterfactual outcomes $\tilde{y}$ nor the treatment assignments $\tilde{t}$ are manually generated. This stands in contrast to existing causal inference benchmarks that must always manually generate $\tilde{y}$ or $\tilde{t}$, even if covariates are based on real data (see, e.g., Louizos et al. (2017)). We find that data generated using our procedure is indeed useful for this task and challenges state-of-the-art estimation methods across both conditional average treatment effect (CATE) and individual treatment effect (ITE)  [^1] It is common to draw no distinction between ITEs and CATEs Vegetabile (2021), but we emphasize these two quantities as distinct: the CATE is the conditional expectation of the ITE, which typically will not explain all ITE variation Lei and Candès (2020). estimation.
1.1 Contribution.
  1. 1.
    We define a procedure for turning any language model and DAG into a sequence-driven structural causal model (SD-SCM). Section˜3 characterizes how an SD-SCM provides access to observational, interventional, and counterfactual distributions over sequential data according to the desired DAG.
  2. 2.
    In Section˜4, we use SD-SCMs to create an example benchmark for causal effect estimation and test a suite of popular estimation methods across CATE and ITE estimation. We find that our benchmark challenges state-of-the-art estimation methods. All benchmark datasets as well as code for SD-SCM-based data generation is available on GitHub. [^2] https://github.com/lbynum/sequence-driven-scms
  3. 3.
    Section˜5 demonstrates how this same technique can underpin auditing language models for (un)desirable causal effects.
Before describing our framework formally, we provide a toy example that illustrates the main points.
Example 1 (Improving your marathon time at the gym).
In this toy example, we use a language model to sample observational and counterfactual data corresponding to two imagined scenarios, each represented by a DAG. The variables we consider will be represented via sets of sequences, where each set can be viewed as a sample space:
  • $\tilde{u}_{1}$ sample space: “My name is xx.” for all $x\in$ {John, Jane, Alice, Bob, Charlie}
  • $\tilde{u}_{2}$ sample space: “I have a friend named xx.” for all $x\in$ {John, Jane, Alice, Bob, Charlie}
  • weather ($\tilde{w}$) sample space: “The weather outside is xx.” for all $x\in$ {sunny, rainy, snowy, cold, hot}
  • gymOrRun ($\tilde{g}$) sample space: “I owe it to myself to go xx.” for $x\in$ {to the gym, for a run outside}
  • marathonTime ($\tilde{m}$) sample space: “After this, my marathon time will xx." for $x\in$ {improve, worsen}
The difference between the two scenarios is what we choose to do with the weather variable $\tilde{w}$. In the first case, we choose DAG $\mathcal{G}_{1}$ where w is a confounder ($\tilde{u}_{1}\rightarrow\tilde{g}\leftarrow\tilde{w}\rightarrow\tilde{m}\leftarrow\tilde{u}_{2}$). In the second case, we choose DAG $\mathcal{G}_{2}$ where $\tilde{w}$ is instead a collider ($\tilde{u}_{1}\rightarrow\tilde{g}\rightarrow\tilde{w}\leftarrow\tilde{m}\leftarrow\tilde{u}_{2}$). Notice we have full control over the DAG we choose, regardless of what we might expect to happen in the real world or be encoded by the language model (where, for example, we would not expect going to the gym to have any impact on the weather). Each of these DAGs, by definition, induces a corresponding factorization of the joint distribution across the 5 variables. The factorization for $\mathcal{G}_{1}$ is $P(\tilde{g}|\tilde{w},\tilde{u}_{1})P(\tilde{m}|\tilde{w},\tilde{u}_{2})P(\tilde{w})P(\tilde{u}_{1})P(\tilde{u}_{2})$ and the factorization for $\mathcal{G}_{2}$ is $P(\tilde{w}|\tilde{g},\tilde{m})P(\tilde{g}|\tilde{u}_{1})P(\tilde{m}|\tilde{u}_{2})P(\tilde{u}_{1})P(\tilde{u}_{2})$.
Simulating an observational study: Our procedure will use a language model to define each of these conditional distributions, instead of defining them manually. In order to observe data that follows the correct structure, we iteratively sample each variable in ancestral order according to the desired DAG, allowing each variable to see only the text of its parents as input. Doing this allows us to use whatever correlations the language model has encoded to define structural equations. For example, to sample a single observation corresponding to $\mathcal{G}_{1}$, the following five phrases (one for each covariate) are sampled using the language model, where [bracketed text] is filled in by querying the model across the corresponding sample space:
  • $\tilde{u}_{1}$ sample: “My name is [Charlie].”
  • $\tilde{u}_{2}$ sample: “I have a friend named [Alice].”
  • $\tilde{w}$ sample: “The weather outside is [cold].”
  • $\tilde{g}|\tilde{u}_{1},\tilde{w}$ sample: “My name is Charlie. The weather outside is cold. I owe it to myself to go [to the gym].”
  • $\tilde{m}|\tilde{u}_{2},\tilde{w}$ sample: “I have a friend named Alice. The weather outside is cold. After this, my marathon time will [improve].”
These five text completions correspond to a single observation, where possible values in each sample space are represented by their index. In other words, we have just observed the data point ($\tilde{u}_{1}$, $\tilde{u}_{2}$, $\tilde{w}$, $\tilde{g}$, $\tilde{m}$) = (4, 2, 3, 0, 0) sampled using DAG $\mathcal{G}_{1}$. Figure˜2 shows the result of repeating this process for 1000 observations with $\mathcal{G}_{1}$ and 1000 observations with $\mathcal{G}_{2}$ using GPT-2 (Radford et al., 2019) as the language model. For $\mathcal{G}_{1}$, we would expect the magnitude of correlation $\rho$ between $P(\tilde{g}=1)$ and $P(\tilde{m}=1)$ to decrease if we condition on confounder $\tilde{w}$. By contrast, for $\mathcal{G}_{2}$, where $\tilde{w}$ is instead a collider, we would expect the magnitude of $\rho$ to instead increase if we condition on $\tilde{w}$. Figure˜2 shows that this is indeed the case — our sampled data reflects the desired causal structure.
Simulating counterfactual data: We can use a similar procedure to simulate interventions instead of observations: we intervene by manually setting an action (in this case, the value of covariate $\tilde{g}$), and we create a counterfactual outcome by additionally setting exogenous variables $\tilde{u}_{1},\tilde{u}_{2}$ and any observed non-descendants of $\tilde{g}$$\{\tilde{w}\}$ for $\mathcal{G}_{1}$ and $\emptyset$ for $\mathcal{G}_{2}$. In Section˜3, we formally define the correspondence of this process to counterfactual versus interventional distributions. This allows us to directly simulate a counterfactual outcome for each of the observed units, choosing $\tilde{g}=1$ or $\tilde{g}=0$ during sampling to generate each unit’s potential outcomes. We can then, for example, test how well a treatment effect estimation method will perform if the estimation method is given only the observational data, i.e., data without any intervention. The right side of Figure˜2 shows prediction error in standard deviation units when using a random forest to predict the sample average treatment effect (SATE) with $P(\tilde{m}=1)$ as the outcome, either using treatment $\tilde{g}$ as the only covariate, or using both $\tilde{g}$ and $\tilde{w}$. As we would expect, including a confounder leads to more accurate effect estimation, while including a collider does not. This demonstrates in a simple way the utility of controlled causal data generation — we can benchmark effect estimation approaches in different settings of interest.
Benchmarking CATE and ITE estimation: Many realistic datasets exist for benchmarking estimation of average treatment effects (ATEs), because ATEs are often feasible to isolate with proper study design. However, there is a lack of such data for benchmarking CATE and ITE estimation, where either the outcomes or treatment assignments must always be manually generated. The key benefit of simulating data this way is that individual-level counterfactual data are observable and controllable. This allows us to not only test ATE estimation methods like in Figure 2, but more importantly to benchmark individual-level effect estimation.
In the remaining sections, we formalize our procedure beyond this toy example and demonstrate how it can be used to generate more complex data that challenges state-of-the-art causal inference methods across both CATE and ITE estimation.
@@ 2 Related work
Curth et al. (2021) lay out four categories of commonly-used methods for semi-synthetic data generation with known causal effects: (1) simulating treatment effects using real baseline outcomes (Knaus et al., 2021); (2) using real covariates but simulating response surfaces (Wendling et al., 2018; Franklin et al., 2014; Hill, 2011); (3) performing biased sampling of randomized data (Gentzel et al., 2021; Dehejia and Wahba, 1999); and (4) constructing (proxies of) counterfactuals and interventions from real or empirical data (Louizos et al., 2017; Gentzel et al., 2019). The paradigm of fitting models to real data and then sampling synthetic data from the fit models is common in many works (Schuler et al., 2017; Neal et al., 2020). In this area, the most closely related works to ours in spirit are those that fit generative models to real datasets such that treatments, outcomes, and covariates — in effect, entirely new datasets — can be sampled, such as Athey et al. (2024) and Neal et al. (2020). While such methods are similar in that they rely on generative models, they are fundamentally different from ours, as they are based on individual datasets that already exist (and already have a fixed causal structure), rather than allowing for arbitrary causal structures to be imagined by a user and then parameterized by a generative model. Our setup is akin to a high-fidelity simulation environment (McDuff et al., 2022) that provides empirical counterfactual data (Gentzel et al., 2019), but without needing to manually design all aspects of the simulation, and in a manner that is instead based on natural language. This work is also loosely related to methods that parameterize structural causal models (SCMs) with generative models or other deep learning components, such as Pawlowski et al. (2020); Sanchez and Tsaftaris (2022), but such methods are geared towards counterfactual inference and learning causal relationships from existing data, rather than flexible data generation.
@@
Our work is not the first to suggest that language models can generate outputs that have casual structure. Many works aim to augment language models with the ability to generate counterfactual data (Chatzi et al., 2024; Li et al., 2023; Betti et al., 2023; Hao et al., 2021; Gat et al., 2023). Counterfactuals and causal reasoning are useful across various natural language processing (NLP) tasks, making this capability of particular interest for ongoing LM research (Wang et al., 2024), and language models with causal reasoning capabilities have a wide variety of applications both within and beyond NLP (Vashishtha et al., 2023; Jin et al., 2023a; Zecevic et al., 2023; Liu et al., 2024; Feder et al., 2021; Kıcıman et al., 2023; Jin et al., 2023b; Gat et al., 2023). We are also not the first to point out that counterfactual data generation with language models is useful for understanding the internal ‘world model’ constructed by an LM and auditing for bias (Fryer et al., 2022). The most similar works to ours that we know of are the contemporaneous works Chatzi et al. (2024) and Ravfogel et al. (2024), which also model counterfactuals in LMs using SCMs. These works focus on how to generate counterfactual strings after network interventions within the LM itself. To achieve this, they leverage the Gumbel-Max trick to infer the noise responsible for generating an input and reuse the same noise (or an inferred noise distribution) to generate a corresponding counterfactual output. Our work is fundamentally different in two key ways. First, we consider semantic interventions rather than network interventions, i.e., modeling causal relationships and counterfactuals all within a semantically meaningful simulation based on a fixed LM. Second, we control the causal structure of the data generation process, taking a DAG as input and generating data according to that DAG.
In more general terms, we focus on how to generate data given a desired causal structure. This capability has important use-cases for downstream tasks like the ones we demonstrate here — generating treatment effects to benchmark effect estimation methods and testing for encoded effects. But more broadly, we provide a generalization of how sequence data and structural causal models can be combined in order to flexibly generate observational, interventional, and counterfactual data for whatever purpose it might be useful.